OpenStudy (anonymous):

Find the exact area of a circle having the given circumference. 3pi

OpenStudy (anonymous):

The formula of circumference is 2piR and the formula of area is pi*R^2 The circumference is C = 3pi 3pi = 2pi*R => R = 3PI/2PI => R= 3/2 A = pi*(3/2)^2 A = 9PI/4